### v1

# Universal quantum control over Majorana zero modes

Abstract
Majorana zero mode (MZM) exhibits inherent resilience to local parametric fluctuations in quantum information processing, due to a topological protection mechanism in the non-Abelian braiding statistics of the anyonic quasiparticles. In this paper, we construct the braiding operations between arbitrary pair of three MZMs under the framework of universal quantum control (UQC). Largely detuned driving fields on the mediator, a local defect of lattice, enable indirect and tunable exchange interaction among arbitrary two MZMs. The braiding operations can then present through the evolution along the universal nonadiabatic passages, whose robustness against the driving-field errors can be substantially enhanced by rapid modulation of the passage-dependent global phase. Moreover, a chiral population transfer on MZMs along the universal passage can be perfectly demonstrated in both clockwise and counterclockwise manners. Our scale-free protocol therefore provides an avenue towards the universal quantum control over MZMs, which is fundamental and essential for topological quantum computation.
I Introduction
Certain problems Bravyi et al. (2018) can show provable advantages of quantum computer over its classical counterpart, i.e., Shor’s algorithm for factoring large numbers Calderbank and Shor (1996); Ekert and Jozsa (1996); Kivlichan et al. (2020) and Grover’s algorithm for amplifying probability of target states Grover (1997); Nagib et al. (2025). However, practical quantum computer demands large-scale quantum systems with extensive gate operations Gidney and Ekerå (2019). Each elementary operation can introduce systematic errors that accumulate exponentially, posing significance of error correction schemes, such as the Shor’s code Shor (1995). That scheme encodes a single logical qubit into nine physical qubits and can detect and correct single-qubit errors. Another paradigm for error correction is topological quantum computation Nayak et al. (2008); Freedman (1998); Kitaev (2001, 2003); Das Sarma et al. (2005); Stern and Halperin (2006); Bonderson et al. (2006), in which qubits are encoded on the non-Abelian anyonic quasiparticles Nayak et al. (2008). Quantum information encoded in those non-local degrees of freedom is believed to maintain intrinsic fault-tolerance against local parametric fluctuations.
Anyons fundamentally bifurcate into the Abelian Wilczek (1990) and non-Abelian categories Read and Green (2000); Ivanov (2001); Alicea et al. (2011); Sarma et al. (2015). The unitary evolutions resulted from braiding Abelian anyons yield the accumulation of phases: the wave function of multiple Abelian anyons acquires a fractional phase factor $e^{i\theta}$, $\theta\neq 0,\pi$, on the exchange of quasiparticles. Bosons ($\theta=0$) and fermions ($\theta=\pi$) are the limiting cases Wilczek (1990). Besides the phase accumulation, the braid group Ivanov (2001); Alicea et al. (2011); Sarma et al. (2015) associated with the non-Abelian anyons enables nontrivial unitary operations among the degenerate many-quasiparticle states. Due to the fact that the qubits encoded in the global degenerate states are robust against the local systematic errors Witzel and Das Sarma (2006), the non-Abelian anyons can serve as the primitive units of topological quantum computation.
Majorana zero mode (MZM), first theoretically predicted in the Kitaev chain model Kitaev (2001), represents the most elementary realization of the non-Abelian anyons. The Kitaev chain models a one-dimensional system of spinless fermions and was once regarded as a somewhat unphysical toy model Elliott and Franz (2015). Nevertheless, the spinless fermions behavior can be simulated through experimentally accessible proxy systems, such as the spin-orbit-coupling electron gases under Zeeman field, as suggested by the subsequent theoretical studies Elliott and Franz (2015); Fu and Kane (2008); Sau et al. (2010); Alicea (2010); Lutchyn et al. (2010); Oreg et al. (2010); Nadj-Perge et al. (2013); Pientka et al. (2013); Gao et al. (2024). Experimental researches for MZM have been conducted in multiple physical platforms, including the superconducting systems Nadj-Perge et al. (2014); Kim et al. (2018); Jiao et al. (2018); Schneider et al. (2020); Yang et al. (2021) and hybrid magnetic-superconductor systems Schneider et al. (2022). Recently, indirect evidences for MZMs have been witnessed in the hybrid semiconductor-superconductor system Quantum (2025), exhibiting a zero-bias conductance peak Das et al. (2012); Deng et al. (2012); Finck et al. (2013); Churchill et al. (2013) and the topological protection against the variations of local parameters Jiao et al. (2018); Yang et al. (2021); Quantum (2025). The braiding of MZMs, however, remains the gold-standard signature for definitively establishing their non-Abelian statistics Dvir et al. (2023).
Conventional braiding operations of MZMs were proposed under the adiabatic condition Ivanov (2001); Stone and Chung (2006); Read (2009); Sau et al. (2011); van Heck et al. (2012); Kraus et al. (2013) and governed by the Berry-Wilczek-Zee holonomy Wilczek and Zee (1984); Cheng et al. (2011). In practice, the implementation of topological quantum computation based on the adiabatic braiding operations is constrained by the extended operation time. First, the prolonged exposure to the environment can induce ruinable decoherence effect on the quantum systems Carmichael (1999). Second, while the exact adiabatic solution is established within infinite evolution time, the realistic operation with a finite clock speed unavoidably entails nonzero non-adiabatic errors Nayak et al. (2008). Such errors will lead to the leakage from the subspace of degenerate ground states to the external Hilbert space. Third, the long operation time of topological quantum computer will increase the possibility of the quasiparticle poisoning Catelani et al. (2011); Karzig et al. (2021); Liu et al. (2024), i.e., a quasiparticle enters the collection of MZMs and changes the parity of the states. Although the measurement-based tools Knapp et al. (2016) have been used to mitigate unwanted transition errors during adiabatic braid operations Bonderson et al. (2008); Bomantara and Gong (2020), a fast and high-fidelity nonadiabatic braiding operation remains desired for MZMs.
Inspired by transitionless quantum driving Berry (2009), the shortcuts to non-Abelian braiding Karzig et al. (2015); Song et al. (2025) exploit the time-dependent and direct exchange interaction between MZMs as the counterdiabatic driving fields. But practical implementation of such interactions in semiconductor-superconductor systems is fundamentally difficult Amundsen et al. (2024). In analog to nonadiabatic holonomic quantum computation Sjöqvist et al. (2012); Liu et al. (2019), the nonadiabatic braiding operations Yu et al. (2025) have been demonstrated between two MZMs mediated by the lattice defect (LD). However, the braiding operations under the parallel-transport condition Jing et al. (2017) is sensitive to the systematic errors of the local driving fields on LD.
We here construct the braiding operations between any pair of three MZMs mediated with LD under local driving fields by universal quantum control (UQC) Jin and Jing (2025a, b, c, d). It is a theoretical framework that unifies the shortcuts to adiabaticity and holonomic quantum transformation and is beyond the parallel-transport condition. The degree of freedom of LD can be eliminated Jin and Jing (2025d); James and Jerke (2007) when the driving fields are largely detuned; and then it enables indirect exchange interaction among MZMs. Consequently, the braiding operations of arbitrary two MZMs can be performed along the universal passages of the system. Our protocol presents fault-tolerant braiding operations that are robust against significant fluctuations in both transition frequency of LD and Rabi frequency of driving field. In addition, the clockwise and counterclockwise population transfers of three MZMs can be demonstrated by the time evolution along the passage that is superposed by LD excited state and MZMs.
The rest part of this paper is structured as follows. In Sec. II.1, we introduce a general model that is consisted of three MZMs coupled to a lattice defect and the full Hamiltonian. In Sec. II.2, we derive the effective Hamiltonian by using the local largely-detuned driving fields on LD, based on which the braiding operations can be established via universal nonadiabatic passages. In Sec II.3, we employ a rapidly-varying global phase to correct systematic errors. Fault-tolerant, fast, and high-fidelity braiding operations can then be constructed for MZMs. Section III devotes to designing the chiral population transfer of three MZMs in both clockwise and counterclockwise manners. The whole work is concluded in Sec. IV. Appendix A presents the UQC theory in the parametric space with ideal control setting, followed by an error-correction extension in Appendix B.
II Theoretical framework
II.1 Model and Hamiltonian
Figure 1: (a) Sketch of an atomic chain system comprising a LD and three one-dimensional NN-site Kitaev chains. LD is coupled to MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$ through local off-resonant driving fields applied to the LD. Inset: the four-level transition diagram for the effective subspace of LD and MZMs. (b) Effective two-level transition diagram for the braiding operation between MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$.
Our braiding protocol is conducted on the atomic-chain system as shown in Fig. 1(a). The system comprises a two-level system (lattice defect) as a mediator, consisting of an excited state $|e\rangle$ and a ground state $|g\rangle$, and three one-dimensional Kitaev chains Kitaev (2001, 2003) of open boundary condition, each of which is composed of $N$ sites. The system Hamiltonian can be written as Kitaev (2001, 2003) ($\hbar\equiv 1$)
$H(t)$ $=\sum_{k=1}^{3}\Big{[}-\mu\sum_{n=1}^{N}c_{n,k}^{\dagger}c_{n,k}-\sum_{n=1}^{N-1}(Jc_{n+1,k}^{\dagger}c_{n,k}+$ (1)
$+gc_{n+1,k}^{\dagger}c_{n,k}^{\dagger}+{\rm H.c.})\Big{]}+\omega(t)\sigma_{+}\sigma_{-},$
where $\sigma_{+}\equiv|e\rangle\langle g|$ ($\sigma_{-}=|g\rangle\langle e|$) and $c_{n,k}^{\dagger}$ ($c_{n,k}$) are the creation (annihilation) operator for LD and the $n$th site of the $k$th Kitaev chain, respectively. The fermions in the Kitaev chains are assumed to have the same transition frequency $\mu$. $\omega(t)$ is the transition frequency of the two-level lattice defect, which is set as a constant in our protocol despite it can be tuned via an external magnetic field Wolfowicz et al. (2021). $J$ and $g$ are the hopping and pairing amplitudes between the fermions in the chains, respectively.
The Hamiltonian in Eq. (1) admits an alternative representation in terms of Majorana fermion quasiparticles, which emerge as bounded pairs in the system. Specifically, the Majorana fermions of the $k$th Kitaev chain, $1\leq k\leq 3$, can be defined as Kitaev (2001, 2003)
$\gamma_{2n-1,k}=c_{n,k}+c_{n,k}^{\dagger},\quad\gamma_{2n,k}=-i(c_{n,k}-c_{n,k}^{\dagger}),\\$ (2)
where $n$ ranges from $1$ to $N$. As defined in Eq. (2), Majorana fermions are characterized by two fundamental properties: (i) self-conjugation, $\gamma_{n,k}^{\dagger}=\gamma_{n,k}$ and (ii) anti-commutation relations $\{\gamma_{n,j},\gamma_{m,k}\}=2\delta_{nm}\delta_{jk}$, where $n$ and $m$ range from $1$ to $2N$. Under the parametric conditions of $\mu=0$ and J=gJ=-g Kitaev (2001, 2003) and Eq. (2), the Hamiltonian (1) is transformed to be
$H(t)=ig\sum_{k=1}^{3}\sum_{n=1}^{N-1}\gamma_{2n+1,k}\gamma_{2n,k}+\omega\sigma_{+}\sigma_{-}.$ (3)
Equations (2) and (3) indicate that there exist two decoupled Majorana zero modes in each Kitaev chain, i.e., $\gamma_{1,k}$ and $\gamma_{2N,k}$.
We are desired to construct the braiding operations among chosen MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, via a mediator LD. In the absence of LD, when the system is initially in the vacuum state $|0\rangle$, the dynamics is constrained in the subspace spanned by the states $|\gamma_{1,k}\rangle=\gamma_{1,k}|0\rangle$. In the presence of the coupled LD under local driving fields, the full Hamiltonian can be expressed as Yu et al. (2025)
$H_{\rm tot}(t)=H(t)+H_{d}(t),$ (4)
with the driving Hamiltonian
$H_{d}(t)=2\sum_{k=1}^{3}\Omega_{k}(t)\cos\left[\int_{0}^{t}\omega_{k}(s)ds+\varphi_{k}\right]|e\rangle\langle\gamma_{1,k}|+{\rm H.c.},$ (5)
where $\Omega_{k}(t)$, $\omega_{k}(t)$, and $\varphi_{k}$ are the Rabi frequency, the driving frequency, and the phase of the driving fields, respectively. While $\gamma_{1,k}$’s involve only a few sites of the Kitaev chains, the driving Hamiltonian in Eq. (5) can be engineered through local pairing and hopping couplings between LD and the proximal sites of the chains. It has been extensively studied, both theoretically Liu et al. (2013); Li et al. (2014); Long et al. (2021); Nathan et al. (2021); Xu et al. (2023) and experimentally Cooper et al. (2019); Miles et al. (2025), in cold atoms and quantum-dot arrays.
Adapt from MZMs, the full Hamiltonian $H_{\rm rot}(t)$ (4) indicates that there are no interactions between LD and the rest Majorana fermions $\gamma_{2n}$ and $\gamma_{2n+1}$, $1\leq n\leq N-1$. Then we can focus on
$H_{\rm tot}(t)$ $=\omega\sigma_{+}\sigma_{-}+\Big{\{}2\sum_{k=1}^{3}\Omega_{k}(t)$ (6)
$\times\cos\left[\int_{0}^{t}\omega_{k}(t^{\prime})dt^{\prime}+\varphi_{k}\right]|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}\Big{\}}.$
In the rotating frame with respect to $H_{0}=\omega\sigma_{+}\sigma_{-}$, it can be rewritten as a time-dependent control Hamiltonian
$H_{I}(t)$ $=\sum_{k=1}^{3}\Omega_{k}(t)\Big{\{}e^{i\int_{0}^{t}[\omega_{k}(s)+\omega]ds+i\varphi_{k}}$ (7)
$+e^{-i\int_{0}^{t}\Delta_{k}(s)ds-i\varphi_{k}}\Big{\}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.},$
where the detuning $\Delta_{k}(t)\equiv\omega_{k}(t)-\omega$. The relevant level transition diagram can be found in the inset of Fig. 1(a).
II.2 Braiding operations via UQC passage
In this subsection, we construct the braiding operations between arbitrary pair of three MZMs, i.e., $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, through a synthesis of the effective Hamiltonian Jin and Jing (2025d); James and Jerke (2007) and UQC theory Jin and Jing (2025a, b, c, d) (see Appendix A for details). In the regime where the mediator LD is under the largely detuned driving fields, the system dynamics is governed by the effective Hamiltonian Jin and Jing (2025d); James and Jerke (2007) as
$H_{\rm eff}(t)=-\frac{i}{2}\left[H_{I}(t),\int_{0}^{t}H_{I}(t_{1})dt_{1}\right].$ (8)
In the control Hamiltonian $H_{I}(t)$ in Eq. (7), the Rabi frequencies are assumed to be isotropic $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$ and the detunings are set as $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ with the scaling detunings $\Delta_{0}$ and $\Delta(t)$. Then under the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l}(t)|,|\Delta_{k}(t)-\Delta_{l}(t)|\gg\Omega_{1}(t)$ with $l\neq j,k\in\{1,2,3\}$ and the near-resonant condition $\Delta_{j}(t)-\Delta_{k}(t)=\Delta(t)\sim\Omega_{1}(t)$, the effective Hamiltonian (8) results in a two-body interaction Hamiltonian:
$H_{\rm eff}(t)=\Omega(t)e^{i\int_{0}^{t}\Delta(s)ds+i\phi_{jk}}|\gamma_{1,j}\rangle\langle\gamma_{1,k}|+{\rm H.c.}$ (9)
with the phase difference $\phi_{jk}=\varphi_{j}-\varphi_{k}$ (assumed to be vanishing for simplicity) and the Rabi frequency
$\Omega(t)=\frac{\Omega_{1}^{2}(t)}{\Delta_{0}+\Delta(t)/2}\approx\frac{\Omega_{1}^{2}(t)}{\Delta_{0}},$ (10)
where the approximation holds under $\Delta_{0}\gg\Delta(t)/2$. In the double-rotated picture with respect to $H_{\rm MZM}=\Delta(t)(|\gamma_{1,k}\rangle\langle\gamma_{1,k}|-|\gamma_{1,j}\rangle\langle\gamma_{1,j}|)/2$, we have
$H_{\rm eff}^{\prime}(t)$ $=\frac{\Delta(t)}{2}(|\gamma_{1,j}\rangle\langle\gamma_{1,j}|-|\gamma_{1,k}\rangle\langle\gamma_{1,k}|)$ (11)
$+\Omega(t)|\gamma_{1,j}\rangle\langle\gamma_{1,k}|+{\rm H.c.},$
whose level transition diagram is shown in Fig. 1(b). $\Delta(t)$ in Eq. (11) is directly controlled by the time-dependent detunings in laboratory, i.e., $\Delta_{j}(t)$ and $\Delta_{k}(t)$, rather than the extra driving fields in the previous protocol for effective Hamiltonian Jin and Jing (2025d).
Using the universal quantum control theory Jin and Jing (2025a, b, c, d) for an effective two-level system, the dynamics of the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ can be described in the ancillary representation spanned by
$|\mu_{1}(\theta,\alpha)\rangle$ $=\cos\theta(t)e^{i\frac{\alpha(t)}{2}}|\gamma_{1,j}\rangle-\sin\theta(t)e^{-i\frac{\alpha(t)}{2}}|\gamma_{1,k}\rangle,$ (12)
$|\mu_{2}(\theta,\alpha)\rangle$ $=\sin\theta(t)e^{i\frac{\alpha(t)}{2}}|\gamma_{1,j}\rangle+\cos\theta(t)e^{-i\frac{\alpha(t)}{2}}|\gamma_{1,k}\rangle,$
where the parameters $\theta(t)$ and $\alpha(t)$ manipulate the populations and the relative phase of the states, respectively. Substituting the ancillary basis states in Eq. (12) to the von Neumann equation (45) with the effective Hamiltonian (11), we have
$\Delta(t)=\dot{\alpha}(t)+2\Omega(t)\cot 2\theta(t)\cos\alpha(t),$ (13)
$\Omega(t)=-\frac{\dot{\theta}(t)}{\sin\alpha(t)}.$
Using Eq. (13), the ancillary basis states in Eq. (12) can be activated as useful nonadiabatic passages. Then due to Eq. (42), the effective evolution operator can be parametrized as
$U(\theta,\alpha)$ $=e^{if(\theta,\alpha)}|\mu_{1}(\theta,\alpha)\rangle\langle\mu_{1}(\theta_{0},\alpha_{0})|$ (14)
$+e^{-if(\theta,\alpha)}|\mu_{2}(\theta,\alpha)\rangle\langle\mu_{2}(\theta_{0},\alpha_{0})|,$
where the subscript 0 indicates the initial boundary condition and the global phase $f(\theta,\alpha)$ is determined by
$\dot{f}(\theta,\alpha)=\Omega(t)\frac{\cos\alpha(t)}{\sin 2\theta(t)}=-\dot{\theta}(t)\frac{\cot\alpha(t)}{\sin 2\theta(t)}.$ (15)
To avoid the singularity of the parameters $\Delta(t)$ and $\Omega(t)$ in Eq. (13), one can take $\theta(t)$, $\alpha(t)$, and $f(\theta,\alpha)$ as independent variables. The conditions in Eq. (13) can then be expressed as
$\Delta(t)=\dot{\alpha}(t)+2\dot{f}(\theta,\alpha)\cos 2\theta(t),$ (16a)
$\dot{\alpha}(t)=-\frac{\ddot{\theta}\dot{f}\sin 2\theta-\ddot{f}\dot{\theta}\sin 2\theta-2\ddot{f}\dot{\theta}^{2}\cos 2\theta}{\dot{f}^{2}\sin^{2}2\theta+\dot{\theta}^{2}},$ (16b)
$\Omega(t)=-\sqrt{\dot{\theta}(t)^{2}+\dot{f}(\theta,\alpha)^{2}\sin^{2}2\theta(t)}.$ (16c)
With Eq. (16b), $f(\theta,\alpha)$ can be simplified as $f(\theta)$.
Figure 2: Fidelity $\mathcal{F}(\theta)$ versus the parameter $\theta(t)$ for the states $|n\rangle$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, during the braiding process with the initial population in MZM (a) $\gamma_{1,j}$ and (b) $\gamma_{1,k}$, respectively. The Kitaev-chain size is $N=100$. $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ where $\Delta(t)$ in Eq. (16a) and it is in the order of $\Delta(t)\sim 10^{-2}\Delta_{0}$. $\Delta_{l}=2\Delta_{0}$. Under the Rabi frequencies of $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$, which satisfy Eqs. (10) and (16c) with $\theta(t)=\pi t/(2T)$ ($T$ is the control period) and $f(\theta,\alpha)=0$. The Rabi frequency is in the order of $\Omega_{1}(t)\sim 10^{-2}\Delta_{0}$.
Equations (12) and (14) indicate that the system can evolve along either $|\mu_{1}(\theta,\alpha)\rangle$ or $|\mu_{2}(\theta,\alpha)\rangle$, depending on the boundary conditions of the parameters $\theta(t)$, $\alpha(t)$, and $f(\theta,\alpha)$. For example, when $\theta(0)=0$, $\theta(T)=\pi/2$ with $T$ the evolution period, and $\alpha(t)=\pi$, the evolution operator in Eq. (14) can become
$U(\theta,\alpha)=-e^{if(\theta,\alpha)}|\gamma_{1,k}\rangle\langle\gamma_{1,j}|+e^{-if(\theta,\alpha)}|\gamma_{1,j}\rangle\langle\gamma_{1,k}|.$ (17)
When $f(\theta,\alpha)=0$, it is further reduced to
$U(\theta,\alpha)=-|\gamma_{1,k}\rangle\langle\gamma_{1,j}|+|\gamma_{1,j}\rangle\langle\gamma_{1,k}|=\exp\left(\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k}\right),$ (18)
which is consistent with the braiding operation between the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ Ivanov (2001); Stone and Chung (2006); Read (2009); Sau et al. (2011); van Heck et al. (2012); Kraus et al. (2013).
Due to the transition diagram in the inset of Fig. 1(a), the performance of our braiding protocol can be evaluated by the fidelities for all the four relevant states: $\mathcal{F}(\theta)\equiv|\langle n|\psi(\theta)\rangle|^{2}$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, where the wave function $|\psi(\theta)\rangle$ is obtained by the Schrödinger equation $i\partial|\psi(\theta)\rangle/\partial t=H_{\rm tot}(t)|\psi(\theta)\rangle$ with the total Hamiltonian $H_{\rm tot}(t)$ (4). Figures 2(a) and (b) present the numerical results of the fidelities starting from different initial MZMs. They are exact the same as the braiding process controlled by $U(\theta,\alpha)$ in Eq. (18), during which the states $|\gamma_{1,l}\rangle$ and $|e\rangle$ are effectively decoupled from the whole evolution. Particularly, in Fig. 2(a), the initial state $|\gamma_{1,j}\rangle$ evolves into an equal superposition of the states $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$ when $\theta=\pi/4$ and then completely transfers to the state $|\gamma_{1,k}\rangle$ when $\theta=\pi/2$. When the system is initially populated on MZM $\gamma_{1,k}$, it can be faithfully transferred to the target MZM $\gamma_{1,j}$ when $\theta(T)=\pi/2$ as shown in Fig. 2(b). In the languages of UQC, Figures 2(a) and (b) simulate the nonadiabatic and transitionless passages $|\mu_{1}\rangle$ and $|\mu_{2}\rangle$ in Eq. (12), respectively.
Our protocol can be straightforwardly generalized from Eq. (7) of $3$ MZMs to the case with $M$ MZMs, i.e., the control Hamiltonian can be scaled to
$H_{I}^{(M)}(t)$ $=\sum_{k=1}^{M}\Omega_{k}(t)\Big{\{}e^{i\int_{0}^{t}[\omega_{k}(s)+\omega]ds+i\varphi_{k}}$ (19)
$+e^{-i\int_{0}^{t}\Delta_{k}(s)ds-i\varphi_{k}}\Big{\}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}.$
The preceding construction of braiding operation can be implemented between any selected pair of MZMs, e.g., the $j$th and $k$th MZMs, $j\neq k$. It is found that under the isotropic condition for Rabi frequencies $\Omega_{l}(t)=\Omega_{1}(t)$ with $1\leq l\leq M$, the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l\neq j,k}(t)|,|\Delta_{k}(t)-\Delta_{l\neq j,k}(t)|\gg\Omega_{1}(t)$, and the near-resonant condition $\Delta_{j}(t)-\Delta_{k}(t)\sim\Omega_{1}(t)$, one can obtain the same two-body effective Hamiltonian in Eq. (9) by Eq. (8). Thus, our braiding protocol is scale free.
II.3 Suppress systematic errors in local defect
While the braiding operations of MZMs remain robust against local perturbations in the other Majorana fermions, this topological protection does not cover the systematic errors about the mediator LD. As indicated by Eq. (47), the unavoidable systematic errors associated with LD can cause undesirable transitions among the universal passages. Incorporating an error correction mechanism in e.g., Eq. (53) to our protocol provides a reliable method to mitigate these systematic errors during the braiding operations of MZMs.
From a pedagogical perspective, the full Hamiltonian $H_{\rm rot}(t)$ in Eq. (4) can be perturbed by: (a) the fluctuation in the splitting frequency of the local defect, i.e.,
$\omega\rightarrow(1+\epsilon)\omega,$ (20)
or (b) the Rabi-frequency fluctuation in the driving field for one of the transitions $|e\rangle\leftrightarrow|\gamma_{1,1}\rangle$, i.e.,
$\Omega_{1}(t)\rightarrow(1+\epsilon)\Omega_{1}(t),$ (21)
where $\epsilon$ is the perturbative coefficient measuring the error magnitude. To suppress the undesirable transitions between the passages by our correction mechanism (53), one can simply set the global phase as
$\frac{\partial f(\theta)}{\partial\theta(t)}=\lambda,$ (22)
where the factor $\lambda$ scales the variation rapidity of $f(\theta)$ relative to $\theta(t)$. $\lambda=0$ represents the parallel-transport condition with a static global phase Jin and Jing (2025c) and $\lambda\gg 1$ describes a rapid-varying global-phase that can suppress the systematic errors in the relevant nonadiabatic passage.
Figure 3: Fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the perturbative coefficient $\epsilon$, controlled by the full Hamiltonian $H_{\rm tot}(t)$ (4) with (a) deviated transition frequency $\omega$ in Eq. (20) and (b) deviated Rabi-frequency $\Omega_{1}(t)$ in Eq. (21), under the global phase $f(\theta)$ in Eq. (22) with various coefficients $\lambda$. The other parameters are the same as Fig. 2(a).
Figures 3(a) and (b) demonstrate the fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the dimensionless error magnitude $\epsilon$, in the presence of the systematic errors in Eqs. (20) and (21), respectively. One can find that the braiding operations incorporating with our correction mechanism ($\lambda>1$) exhibits superior performances in comparison to those with no correction mechanism ($\lambda=0$). In Fig. 3(a) for $\lambda=0$, the fidelity decreases to $\mathcal{F}=0.924$ when $\epsilon=-0.05$ and $\mathcal{F}=0.807$ when $\epsilon=0.05$. Under our correction mechanism, the fidelity admits a lower bound as $\mathcal{F}>0.963$ when $\lambda=2$ and $\mathcal{F}>0.995$ when $\lambda\geq 5$, in the whole range of $\epsilon\in[-0.05,0.05]$. As shown in Fig. 3(b), the braiding operation with our correction mechanism maintains a stronger robustness against the deviation in the Rabi-frequency of a local driving field. In particular, when $\lambda=0$, the fidelity is about $\mathcal{F}=0.962$ when $\epsilon=-0.1$ and $\mathcal{F}=0.968$ when $\epsilon=0.1$. The adverse effects from the local errors can be suppressed so that $\mathcal{F}>0.980$ when $\lambda=2$, and $\mathcal{F}>0.998$ when $\lambda\geq 5$.
III Chiral population transfer among Majorana zero modes
Figure 4: Four-level transition diagram with the coupling between LD and MZMs $\gamma_{1,k}$, $k=1,2,3$. Black solid lines describe the off-resonant driving fields on the transition $|e\rangle\leftrightarrow|\gamma_{1,k}\rangle$ with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The grey dashed lines refer to the population transfer among MZMs along a clockwise or counterclockwise direction.
Beyond the large-detuning regime, we can apply the universal quantum control theory Jin and Jing (2025a) (see Appendix A for details) to realize a chiral population transfer Wang et al. (2016); Liu et al. (2020); Qi and Jing (2022) among the three MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$. As shown in Fig. 4, a LD coupled to MZM $\gamma_{1,k}$’s, $k=1,2,3$, by the local driving field with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The full Hamiltonian can be obtained by Eq. (7), which is recalled as below
$H_{I}(t)$ $=\sum_{k=1}^{3}\Omega_{k}(t)\Big{\{}e^{i\int_{0}^{t}[\omega_{k}(s)+\omega]ds+i\varphi_{k}}$ (23)
$+e^{-i\int_{0}^{t}[\omega_{k}(s)-\omega]ds-i\varphi_{k}}\Big{\}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}.$
Formally the driving frequencies satisfy $\omega_{k}(t)=-\omega+\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$. In the large transition-frequency regime $\omega\gg\Omega_{k}(t),\Delta_{e}(t),\tilde{\Delta}_{k}(t)$, $1\leq k\leq 3$, Eq. (23) can be reduced to
$H_{I}(t)=\sum_{k=1}^{3}\Omega_{k}(t)e^{i\int_{0}^{t}[-\tilde{\Delta}_{k}(s)+\Delta_{e}(s)]ds+i\varphi_{k}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.},$ (24)
under the rotating-wave approximation. In the double-rotated frame with respect to $H_{0}=-\Delta_{e}(t)|e\rangle\langle e|-\sum_{k=1}^{3}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$, the Hamiltonian (24) can be transformed as
$H_{I}(t)$ $=\Delta_{e}(t)|e\rangle\langle e|+\sum_{k=1}^{3}\Big{\{}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$ (25)
$+\left[\Omega_{k}(t)e^{i\varphi_{k}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}\right]\Big{\}}.$
According to UQC recipe for two-band systems Jin and Jing (2025b), the system dynamics can be described in the ancillary picture spanned by four ancillary basis states, which can be chosen as
$|\mu_{1}(t)\rangle$ $=\cos\theta(t)|\gamma_{1,1}\rangle-\sin\theta(t)e^{-i\alpha_{1}(t)}|\gamma_{1,2}\rangle,$ (26)
$|\mu_{2}(t)\rangle$ $=\cos\phi(t)|b_{1}(t)\rangle-\sin\phi(t)e^{-i\alpha_{2}(t)}|\gamma_{1,3}\rangle,$
$|\mu_{3}(t)\rangle$ $=\cos\chi(t)|b_{2}(t)\rangle-\sin\phi(t)e^{-i\alpha_{3}(t)}|e\rangle,$
$|\mu_{4}(t)\rangle$ $=\sin\chi(t)|b_{2}(t)\rangle+\cos\phi(t)e^{-i\alpha_{3}(t)}|e\rangle.$
Here the bright states are given by
$|b_{1}(t)\rangle$ $=\sin\theta(t)|\gamma_{1,1}\rangle+\cos\theta(t)e^{-i\alpha_{1}(t)}|\gamma_{1,2}\rangle,$ (27)
$|b_{2}(t)\rangle$ $=\sin\phi(t)|b_{1}(t)\rangle+\cos\phi(t)e^{-i\alpha_{2}(t)}|\gamma_{1,3}\rangle,$
which maintain the orthogonality with the relevant ancillary basis states, i.e., $\langle\mu_{n}(t)|b_{n}(t)\rangle=0$, $n=1,2$. The time-dependent parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$ control the state populations, whereas $\alpha_{1}(t)$, $\alpha_{2}(t)$, and $\alpha_{3}(t)$ determine the relative phases among states. Equation (26) constitute a completed and orthornormal basis for the Hilbert space spanned by LD excited state and three MZMs.
Although each basis state in Eq. (26) serves as a nonadiabatic passage when it satisfies the von Neumann equation (45) with the Hamiltonian $H(t)$ (25), it is found that only $|\mu_{3}(t)\rangle$ and $|\mu_{4}(t)\rangle$ can be used to realize the chiral transfer of MZMs. Without loss of generality, we substitute $|\mu_{4}(t)\rangle$ into the von Neumann equation (45) and then obtain the conditions for the Rabi-frequencies
$\Omega_{1}(t)$ $=-\Big{(}\dot{\chi}\sin\phi\sin\theta+\dot{\phi}\tan\chi\cos\phi\sin\theta+\dot{\theta}\tan\chi$ (28)
$\times\sin\phi\cos\theta\Big{)}/\sin(\varphi_{1}+\alpha_{2}),$
$\Omega_{2}(t)$ $=-\Big{(}\dot{\chi}\sin\phi\cos\theta+\dot{\phi}\tan\chi\cos\phi\cos\theta-\dot{\theta}\tan\chi$
$\times\sin\phi\sin\theta\Big{)}/\sin(\varphi_{2}-\alpha_{0}+\alpha_{2}),$
$\Omega_{3}(t)$ $=-\left(\dot{\chi}\cos\phi-\dot{\phi}\tan\chi\sin\phi\right)/\sin(\varphi_{3}-\alpha_{1}+\alpha_{2}),$
and those for the detunings
$\tilde{\Delta}_{1}(t)=-\Omega_{1}(t)\frac{\cot\chi}{\sin\phi\sin\theta}\cos(\varphi_{1}+\alpha_{2}),$ (29)
$\tilde{\Delta}_{2}(t)=-\Omega_{2}(t)\frac{\cot\chi}{\sin\phi\cos\theta}\cos(\varphi_{2}-\alpha_{0}+\alpha_{2})$
$+\dot{\alpha}_{0}\sin\chi\sin\phi\cos\theta,$
$\tilde{\Delta}_{3}(t)=-\Omega_{3}\frac{\cot\chi}{\cos\phi}\cos(\varphi_{3}-\alpha_{1}+\alpha_{2})+\dot{\alpha}\sin\chi\cos\phi,$
$\Delta_{e}(t)=\dot{\alpha}_{2}-\Omega_{1}(t)\sin\chi\sin\phi\sin\theta\cos(\varphi_{1}+\alpha_{2})$
$-\Omega_{2}(t)\sin\chi\sin\phi\cos\theta\cos(\varphi_{2}-\alpha_{0}+\alpha_{2})-\Omega_{3}\sin\chi$
$\times\cos\phi\cos(\varphi_{3}-\alpha_{1}+\alpha_{2}).$
With a appropriate choice of the parameters of $|\mu_{4}(t)\rangle$ in Eq. (26), the parameters in the lab can be reversely engineered by Eqs. (28) and (LABEL:CondDetu), then the system evolution can be precisely controlled to follow the desired passage. A complete loop of chiral population transfer of the three MZMs can be consisted of three sequential stages, each lasing identical time interval $T$. In particular, for the system initialized in the state $|\gamma_{1,1}\rangle$, the clockwise population transfer proceeds through Stage (i) $|\gamma_{1,1}\rangle\rightarrow|\gamma_{1,2}\rangle$ when $t\in[0,T]$, Stage (ii) $|\gamma_{1,2}\rangle\rightarrow|\gamma_{1,3}\rangle$ when $t\in[T,2T]$, and Stage (iii) $|\gamma_{1,3}\rangle\rightarrow|\gamma_{1,1}\rangle$ when $t\in[2T,3T]$. They are determined by the boundary conditions for the parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$: $\theta(0)=\phi(0)=\chi(0)=\pi/2$, $\theta(T)=\pi$, $\phi(T)=\chi(T)=\pi/2$, $\phi(2T)=0$, $\chi(2T)=\pi/2$, and $\theta(3T)=\phi(3T)=\chi(3T)=\pi/2$. Without loss of generality, $\theta(t)$ and $\phi(t)$ for Stages (i), (ii), and (iii) of the $k$th loop, $k\geq 1$, can be set as
$\theta(t)$ $=\Phi(t)+\frac{\pi}{2},\quad\phi(t)=2\Phi(t)+\frac{\pi}{2},$ (30)
$\theta(t)$ $=\Phi(t)+\frac{\pi}{2},\quad\phi(t)=\Phi(t),$
$\theta(t)$ $=\phi(t)=\Phi(t),$
respectively, with $\Phi(t)\equiv\pi[t-3(k-1)T]/(2T)$, and $\chi(t)$ for the whole loop can be set as
$\chi(t)=\frac{\pi}{2}\left[1+\frac{\sin{(\frac{\pi t}{T})}}{1+\sin^{2}(\frac{\pi t}{T})}\right].$ (31)
Figure 5: Fidelity dynamics $\mathcal{F}_{n}(t)$ about various states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, for two loops of (a) clockwise and (b) counterclockwise population transfers. The Rabi frequencies and the detunings of the Hamiltonian (25) are set by Eqs. (28) and (LABEL:CondDetu) with $\varphi_{1}-\alpha_{2}=\pi/2$, $\varphi_{2}-\alpha_{0}+\alpha_{2}=\pi/2$, $\varphi_{3}-\alpha_{1}+\alpha_{2}=\pi/2$ and (a) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (30) and (31) and (b) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (32), (33), and (31). The Rabi frequencies are in the same order as those in Fig. 2.
Similarly, the counterclockwise population transfer proceeds through Stage (i) $|\gamma_{1,1}\rangle\rightarrow|\gamma_{1,3}\rangle$, Stage (ii) $|\gamma_{1,3}\rangle\rightarrow|\gamma_{1,2}\rangle$, and Stage (iii) $|\gamma_{1,2}\rangle\rightarrow|\gamma_{1,1}\rangle$. For the $k$th loop, $\theta(t)$ for Stage (i) and Stage (ii-iii) can be set as
$\theta(t)=\Phi(t)+\frac{\pi}{2},\theta(t)=\Phi(t)+\pi,$ (32)
respectively, and $\phi(t)$ for Stages (i-ii) and Stage (iii) can be chosen as
$\phi(t)=\Phi(t)+\frac{\pi}{2},\quad\phi(t)=\Phi(t),$ (33)
respectively. Both $\chi(t)$ and $\Phi(t)$ hold the preceding definition.
Figure 5 demonstrates the fidelity dynamics $\mathcal{F}_{n}(t)$ about the four relevant states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, during the chiral population transfer. It is found that the population can propagate with a unit fidelity in both clockwise and counterclockwise manners. In Fig. 5(a) for the clockwise population transfer, the initial population on the first MZM $\gamma_{1,1}$ can be completely transferred to the second one $\gamma_{1,2}$ when $t=T$, despite the unwanted states are temporally populated during Stage (i). We have up to $\mathcal{F}_{e}\approx\mathcal{F}_{\gamma_{1,3}}\approx 0.5$ when $t=0.5T$. At the end of Stage (ii), the population transfers completely to the third MZM $\gamma_{1,3}$ at $t=2T$. And the loop is completed by the population returning to the state $|\gamma_{1,1}\rangle$ at $t=3T$. During the second loop $t\in[3T,6T]$, we have another perfect chiral transfer in the clockwise manner. Figure 5(b) demonstrates the counterclockwise population transfer with the sequences: the population in the state $|\gamma_{1,1}\rangle$ can be transferred to the state $|\gamma_{1,3}\rangle$ when $t=T$, followed by the state $|\gamma_{1,2}\rangle$ when $t=2T$, and then back to the state $|\gamma_{1,1}\rangle$ when $t=3T$. The second loop of the counterclockwise population transfer exhibits identical behavior of the first loop.
In conventional protocols for chiral transfer, both spin chirality Liu et al. (2020) and bosonic-mode chirality Qi and Jing (2022) depend on an effective time-reversal-symmetry broken Hamiltonian, from either three-body interaction or Floquet engineering. Normally the Hamiltonian is obtained by the adiabatic elimination over ancillary state or mode in the large-detuning regime, resulting in a comparatively weak effective coupling strength between interested systems. Thus, these protocols are susceptible to the environment noise due to the long running period. Moveover, they demand a precise control over the local phases of individual driving fields. In sharp contrast, our protocol enables the chiral population transfer in a nonadiabatic way. It does not invoke synthesizing an artificial magnetic field and does not require particular $\varphi_{n}$ for each driving field.
Our protocol of chiral population transfer can also be scaled to the case of $M$ MZMs described by Eq. (19). Particularly, by setting the driving frequencies as $\omega_{k}(t)=-\omega+\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$ and the transition frequency $\omega\gg\Omega_{k}(t),\Delta_{e}(t),\tilde{\Delta}_{k}(t)$ for $1\leq k\leq M$, the Hamiltonian (19) can become
$H_{I}^{(M)}(t)$ $=\Delta_{e}(t)|e\rangle\langle e|+\sum_{k=1}^{M}\Big{\{}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$ (34)
$+\left[\Omega_{k}(t)e^{i\varphi_{k}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}\right]\Big{\}}$
in the rotating frame with respect to $H_{0}=-\Delta_{e}(t)|e\rangle\langle e|-\sum_{k=1}^{M}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$, under the rotating-wave approximation. In this $1+M$-dimensional two-band model, the recipe Jin and Jing (2025b) provides two ancillary basis states, as the superpositions of the LD excited state and MZMs, either of which can be activated as a useful passage for chiral population transfer.
IV Conclusion
In the framework of universal quantum control theory, we construct robust, fast, and high-fidelity braiding operations for Majorana zero modes that are coupled to a common two-level local defect. Also, we propose a protocol to realize the perfect chiral population transfer among three MZMs. The braiding operation between an arbitrary pair of MZMs is based on the mediated exchange interaction by using the largely detuned driving fields at LD and the subsequent nonadiabatic passage. Moreover, our universal passage could employ a rapidly-varying global phase instead of extra driving fields to significantly suppress the systematic errors in LD and the driving fields on it. In the large splitting regime of the local defect, we can construct an effective Hamiltonian for the excited state of LD and the three MZMs, by which the population can propagate in both clockwise and counterclockwise manners among three MZMs with a unit fidelity. Our protocol demonstrates fault-tolerant, nonadiabatic and scalable braiding operations and the chiral population transfer with no time-reversal-symmetry broken Hamiltonian, which are fundamental and essential for the topological quantum computation and quantum information.
Acknowledgments
We acknowledge grant support from the Science and Technology Program of Zhejiang Province (No. 2025C01028).
Appendix A UQC theory in parametric space
This appendix extends the universal quantum control theory Jin and Jing (2025a, b, c, d) from the time domain to the parametric space. Consider a $K$-dimensional system under a general effective Hamiltonian $H_{\rm eff}(\vec{R})$ with a varying control parameter vector $\vec{R}\equiv(R_{1},R_{2},\cdots,R_{M})$. $\dot{\vec{R}}\equiv d\vec{R}/dt=0$ means adiabatic condition. $H_{\rm eff}(\vec{R})$ can be obtained by such as Eq. (8), exhibiting an explicit or implicit time dependence, i.e., $R_{j}\equiv R_{j}(t)$ with $1\leq j\leq M$. The system dynamics is described by the parameter-dependent Schrödinger equation
$i\frac{d|\psi_{n}(\vec{R})\rangle}{dt}=H_{\rm eff}(\vec{R})|\psi_{n}(\vec{R})\rangle,$ (35)
where $|\psi_{n}(\vec{R})\rangle$’s are the pure-state solutions, $1\leq n\leq K$, constituting a completed and orthonormal set of the interested system.
In UQC theory Jin and Jing (2025a, b, c), it is convenient to treat the system dynamics within an ancillary picture spanned by a completed and orthonormal set of ancillary basis states, i.e., $|\mu_{k}(\vec{R})\rangle$’s, $1\leq k\leq K$. $|\mu_{k}(\vec{R})\rangle$’s span the same Hilbert space as $|\psi_{n}(\vec{R})\rangle$’s. In the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, where $\vec{R}_{0}$ is a fixed parameter vector representing but not limited to the boundary condition $\vec{R}(t=t_{0})$, the Hamiltonian can be transformed as
$H_{\rm rot}(\vec{R})=V^{\dagger}(\vec{R})H_{\rm eff}(\vec{R})V(\vec{R})-iV^{\dagger}(\vec{R})\dot{V}(\vec{R})$ (36)
$=\sum_{k=1}^{K}\sum_{n=1}^{K}\Big{[}\langle\mu_{k}(\vec{R})|H_{\rm eff}(\vec{R})|\mu_{n}(\vec{R})\rangle$
$-i\sum_{j=1}^{M}\dot{R}_{j}\langle\mu_{k}(\vec{R})|\partial_{R_{j}}\mu_{n}(\vec{R})\rangle\Big{]}|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|$
$=\sum_{k=1}^{K}\sum_{n=1}^{K}\Big{[}\langle\mu_{k}(\vec{R})|H_{\rm eff}(\vec{R})|\mu_{n}(\vec{R})\rangle$
$-i\dot{\vec{R}}\cdot\langle\mu_{k}(\vec{R})|\nabla_{\vec{R}}|\mu_{n}(\vec{R})\rangle\Big{]}|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|$
$\equiv\sum_{k=1}^{K}\sum_{n=1}^{K}\left[\mathcal{H}_{kn}(\vec{R})-\mathcal{A}_{kn}(\vec{R})\right]|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|$
$=V^{\dagger}(\vec{R})\left[\mathcal{H}(\vec{R})-\mathcal{A}(\vec{R})\right]V(\vec{R}),$
where the dynamical component $\mathcal{H}_{kn}(\vec{R})$ and the gauge-potential component $\mathcal{A}_{kn}(\vec{R})$ Kolodrubetz et al. (2017); Claeys et al. (2019); Takahashi and del Campo (2024) constitute the elements of the matrixes $\mathcal{H}(\vec{R})$ and $\mathcal{A}(\vec{R})$, respectively, at the $k$th row and the $n$th column. The purely geometrical matrix $\mathcal{A}(\vec{R})$ relates uniquely to the manifold geometry determined by the ancillary basis states $|\mu_{k}(\vec{R})\rangle$’s. The Schrödinger equation (35) can then be transformed as
$i\frac{d|\psi_{n}(\vec{R})\rangle_{\rm rot}}{dt}=H_{\rm rot}(\vec{R})|\psi_{n}(\vec{R})\rangle_{\rm rot}$ (37)
with the rotated pure-state solutions $|\psi_{n}(\vec{R})\rangle_{\rm rot}$
$|\psi_{n}(\vec{R})\rangle_{\rm rot}\equiv V^{\dagger}(\vec{R})|\psi_{n}(\vec{R})\rangle.$ (38)
The parameter-dependent Schrödinger equation (37) can be exactly solved when $H_{\rm rot}(\vec{R})$ is diagonalizable Jin and Jing (2025a, b, c), particularly, when Eq. (36) can be simplified as
$H_{\rm rot}(\vec{R})=\sum_{k=1}^{K}\left[\mathcal{H}_{kk}(\vec{R})-\mathcal{A}_{kk}(\vec{R})\right]|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{k}(\vec{R}_{0})|.$ (39)
In this case, the evolution operator $U_{\rm rot}(\vec{R})$ can be directly obtained by Eq. (39) as
$U_{\rm rot}(\vec{R})=\sum_{k=1}^{K}e^{if_{k}(\vec{R})}|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{k}(\vec{R}_{0})|,$ (40)
where the global phases $f_{k}(\vec{R})$’s are defined as
$f_{k}(\vec{R})\equiv\int_{\vec{R}_{0}}^{\vec{R}}\frac{1}{\dot{\vec{R}}^{\prime}}\left[\mathcal{A}_{kk}(\vec{R}^{\prime})-\mathcal{H}_{kk}(\vec{R}^{\prime})\right]d\vec{R}^{\prime}.$ (41)
Transformed back to the original picture via Eq. (38), the evolution operator in the original picture can be written as
$U_{0}(\vec{R})=V(\vec{R})U_{\rm rot}(\vec{R})=\sum_{k=1}^{K}e^{if_{k}(\vec{R})}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|.$ (42)
Due to the theory of UQC, the diagonalization condition of $H_{\rm rot}(\vec{R})$ (36) in the $k$th ancillary basis state $|\mu_{k}(\vec{R})\rangle$ is given by
$\left[H_{\rm eff}(\vec{R})-\mathcal{A}(\vec{R}),\Pi_{k}(\vec{R})\right]=0$ (43)
with the projection operator $\Pi_{k}(\vec{R})\equiv|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|$. It admits a compact form as
$\left[H_{\rm eff}(\vec{R}),\Pi_{k}(\vec{R})\right]=\left[\mathcal{A}(\vec{R}),\Pi_{k}(\vec{R})\right]$ (44)
$=$ $i\sum_{j=1}^{M}\dot{R}_{j}\left[|\partial_{R_{j}}\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|+|\mu_{k}(\vec{R})\rangle\langle\partial_{R_{j}}\mu_{k}(\vec{R})|\right]$
$=$ $i\dot{\vec{R}}\cdot\Big{[}\nabla_{\vec{R}}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R})|+|\mu_{k}(\vec{R})\rangle\langle\mu_{k}{(\vec{R})}|\nabla_{\vec{R}}\Big{]}$
$=$ $i\dot{\vec{R}}\cdot\nabla_{\vec{R}}\Pi_{k}(\vec{R}),$
where the second equivalence used the relation for the gauge field, i.e., $\mathcal{A}(\vec{R})|\mu_{k}(\vec{R})\rangle=i\sum_{j=1}^{M}\dot{R}_{j}|\partial_{R_{j}}\mu_{k}(\vec{R})\rangle$, as defined in Eq. (36). Equation (44) can be recast as a von Neumann-type equation Jin and Jing (2025a, b, c):
$\frac{d}{dt}\Pi_{k}(\vec{R})=-i\left[H_{\rm eff}(\vec{R}),\Pi_{k}(\vec{R})\right].$ (45)
When $k$ in Eq. (45) runs from 1 to $K$, $H_{\rm rot}(\vec{R})$ can be completely diagonalized with a full rank.
Appendix B UQC theory in parametric space with error correction
This appendix provides a recipe for the universal quantum control theory in parametric space to suppress the systematic errors Jin and Jing (2025c). When unavoidable fluctuations appear in the control parameters, the system dynamics is actually subject to the full Hamiltonian $H(\vec{R})=H_{\rm eff}(\vec{R})+\epsilon H_{1}(\vec{R})$. Therefore, the parameter-dependent Schrödinger equation (35) becomes
$i\frac{d|\psi_{n}(\vec{R})\rangle}{dt}=H(\vec{R})|\psi_{n}(\vec{R})\rangle,$ (46)
where $H_{1}(\vec{R})$ is the error Hamiltonian and $\epsilon$ is a perturbative coefficient.
Consider the system dynamics in the rotating frame with respect to $V(\vec{R})\equiv\sum_{k=1}^{K}|\mu_{k}(\vec{R})\rangle\langle\mu_{k}(\vec{R}_{0})|$, the ideal rotated Hamiltonian in Eq. (39) can be perturbed by the error Hamiltonian as
$H_{\rm rot}(\vec{R})$ $=-\sum_{k=1}^{K}\dot{\vec{R}}\cdot\nabla_{\vec{R}}f_{k}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{k}(\vec{R}_{0})|$ (47)
$+\epsilon\sum_{k=1}^{K}\sum_{n=1}^{K}\mathcal{D}_{kn}^{\rm err}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|,$
where $\mathcal{D}_{kn}^{\rm err}(\vec{R})\equiv\langle\mu_{k}(\vec{R})|H_{1}(\vec{R})|\mu_{n}(\vec{R})\rangle$. Equation (47) indicates that the errors can yield unwanted transitions of the system from the passage $|\mu_{k}(\vec{R})\rangle$ to the passage $|\mu_{n}(\vec{R})\rangle$, $n\neq k$.
To identify and cancel the adverse effects induced by systematic errors, the system dynamics can be considered in the double-rotated picture with respect to $U_{\rm rot}(\vec{R})$ in Eq. (40). In particular, we have
$H_{\rm err}(\vec{R})$ $=U_{\rm rot}^{\dagger}(\vec{R})H_{\rm rot}(\vec{R})U_{\rm rot}(\vec{R})-iU_{\rm rot}^{\dagger}(\vec{R})\dot{U}_{\rm rot}(\vec{R})$ (48)
$=\epsilon\sum_{k=1}^{K}\sum_{n=1}^{K}\tilde{\mathcal{D}}_{kn}^{\rm err}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle\langle\mu_{n}(\vec{R}_{0})|$
where
$\tilde{\mathcal{D}}_{kn}^{\rm err}(\vec{R})\equiv\langle\mu_{k}(\vec{R})|H_{1}(\vec{R})|\mu_{n}(\vec{R})\rangle e^{-i[f_{k}(\vec{R})-f_{n}(\vec{R})]}.$ (49)
Using the Magnus expansion Blanes et al. (2009), the evolution operator for $H_{\rm err}(\vec{R})$ in Eq. (48) can be expanded as (see Ref. Jin and Jing (2025c) for details)
$U_{\rm err}(\vec{R})$ (50)
$=1-i\int_{\vec{R}_{0}}^{\vec{R}}d\vec{R}^{\prime}\frac{H_{\rm err}(\vec{R}^{\prime})}{\dot{\vec{R}}^{\prime}}-\frac{1}{2}\Big{\{}\left[\int_{\vec{R}_{0}}^{\vec{R}}d\vec{R}^{\prime}\frac{H_{\rm err}(\vec{R}^{\prime})}{\dot{\vec{R}}^{\prime}}\right]^{2}$
$+\int_{\vec{R}_{0}}^{\vec{R}}d\vec{R}^{\prime}\left[\frac{H_{\rm err}(\vec{R}^{\prime})}{\dot{\vec{R}}^{\prime}},\int_{\vec{R}_{0}}^{\vec{R}^{\prime}}d\vec{R}^{\prime\prime}\frac{H_{\rm err}(\vec{R}^{\prime\prime})}{\dot{\vec{R}}^{\prime\prime}}\right]\Big{\}}+\cdots.$
To evaluate the impact of the systematic errors, one can define the overlap or fidelity between the instantaneous state governed by $U_{\rm err}(\vec{R})$ in Eq. (50) and the target passage $|\mu_{k}(\vec{R}_{0})\rangle$ in the rotating frame Jin and Jing (2025c) as
$\mathcal{F}$ $\equiv|\langle\mu_{k}(\vec{R}_{0})|U_{\rm err}(\vec{R})|\mu_{k}(\vec{R}_{0})\rangle|^{2}$ (51)
$\approx 1-\epsilon^{2}\sum_{n=1,n\neq k}^{K}\left|\int_{\vec{R}_{0}}^{\vec{R}}\frac{\tilde{\mathcal{D}}_{kn}^{\rm err}(\vec{R}^{\prime})}{\dot{\vec{R}}^{\prime}}d\vec{R}^{\prime}\right|^{2}+\mathcal{O}(\epsilon^{3}).$
The error up to the second-order can be cancelled under the error correction mechanism Jin and Jing (2025c)
$\left|\nabla_{\vec{R}}\left[f_{k}(\vec{R})-f_{n}(\vec{R})\right]\right|\gg\left|\nabla_{\vec{R}}\left[\langle\mu_{k}(\vec{R})|H_{1}(\vec{R})|\mu_{n}(\vec{R})\rangle\right]\right|.$ (52)
This result is supported by the nonperturbative dynamical decoupling method Jing et al. (2013, 2015) and the Riemann-Lebesgue Lemma in mathematics.
When the control parameter vector $\vec{R}$ only involves one variable $R$, i.e., $\vec{R}\rightarrow R$, Eq. (52) can be reduced to
$\left|\frac{\partial}{\partial R}[f_{k}(R)-f_{n}(R)]\right|\gg\left|\frac{\partial\langle\mu_{k}(R)|H_{1}(R)|\mu_{n}(R)\rangle}{\partial R}\right|.$ (53)
Also, the gradient inequality with respect to the control parameters in Eq. (52) can be reformulated as a gradient inequality with respect to time:
$\left|\frac{\partial}{\partial t}[f_{k}(t)-f_{n}(t)]\right|\gg\frac{\partial}{\partial t}\left[\langle\mu_{k}(t)|H_{1}(t)|\mu_{n}(t)\rangle\right],$ (54)
which is consistent with that in Ref. Jin and Jing (2025c).